Questions 1: A firm has the marginal private costs MPC = 2 + q. The aggregate demand for q is AD = 10 -q. What is the market equilibrium for q?

Question 2: Producing a unit of q has the marginal external costs of 2 in addition to the private costs described in question 1. For the demand function given in question 1, what is the socially efficient q?

Question 3: Assume that protecting a threatened species would produce non-use benefits of 1 million per year. For a discount rate of 5 %, what is the present value of these benefits?

Question 4: Assume that protecting the species would reduce land values by 10 millions because of use restriction (land conversion etc.). How much does the protection reduce the annual (yearly) profits from using the land if the interest rate is 5 % and there is a constant stream of benefits from using the land?
Answer #1

Q1) at market eqm

Market demand is Marginal Benefit curve

So at eqm, AD = PMC

10-q = 2+q

8 = 2q

q* = 4

Q2) MEC = 2

SMC = sovial marginal cost

SMC = PMC + MEC

= 2+q + 2

= 4+q

At eqm, MB = SMC

10-q = 4+q

6 = 2q

q" = 3, socially efficient

C) Present value = 1/(1+ .05)

= 1/1.05

= .9524 $ million

its mandatory to answer only first question